OpenStudy (clamin):

PLEASE HELP!! At how many points does the graph of the function below intersect the x-axis? y=4x^2-6x+1 a. 0 b. 1 c. 2 please show me how to solve it

OpenStudy (anonymous):

When the function intersects the x-axis, y has to be 0. \[y=0\] So you insert that, and solve the equation: \[4x^2-6x+1=0\]

OpenStudy (anonymous):

The general formula for solving x here is: \[\frac{ -b \pm \sqrt{D} }{ 2a }\], where \[D=b^2-4ac\]

OpenStudy (anonymous):

And actually, since you only need to find how many solutions there is, then you just need to solve D. If D is greater than 0, then there is two solutions, if it is equal to 0 there is one solution, and if its below 0 there is zero solutions.